• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gron

    Log entry time 08:02:01 on April 03, 2012

    Entry number 369512

    keyword=Shift Summary - OWL Shift

    Shift Summary: Apr 3rd, OWL

    Shift Crew: SL: Guy Ron

    TO: Toby Badman

    3rd: Paul Gueye

    12:00 - Started shift, RHRS at 1.078 GeV/c (see previous HALOG entry). LHRS 1.0028 GeV/c.

    1:35 - Finished the 1.003 GeV/c setting (with data just in the left arm, right arm stuck at the 1.078 momentum).

    2:00 - Problem with an FZ magnet, MCC called in an expert and will go down into the hall to try to fix it. Moving to controlled access.

    2:00 - IHWP moved out by Hall C.

    2:40 - Still waiting on MCC to fix the FZ magnet problem.

    4:04 - Lumi 3 tripping. Trying to ramp it up slowly but it still trips. Need to fix it tomorrow during beam studies.

    4:04 - MCC expert replacing the regulation module on the FZ magnet.

    4:25 - MCC expert has replaced the regulation module but this has not fixed the problem. Currently they are replacing the control board.

    4:47 - MCC giving us beam back after replacing the control board on the Danphysik PS. Getting some tune beam.

    5:57 - MCC still having problems tuning the beam to our requested position.

    6:25 - Put the carbon hole target in to try to see the beam.

    6:35 - Although the hole doesn't look great (maybe offset by about ~1mm) we're starting dilution runs so as not to waste too much time, since we have beam studies coming up.

    7:15 Continue taking data with the current kinematic until 8am when we go down for beam studies.

    8:00 Shift ends. Next shift with data should move to the next kinematic 0.933GeV/c.